OpenStudy (anonymous):

Factor completely. x^3 + 6x^2 - 4x - 24 A.) (x + 6)(x - 2)(x + 2) B.) (x + 2)(x - 6)(x + 2) C.) (x - 6)(x + 2)(x + 2)

OpenStudy (anonymous):

x - 2 is a factor because f(2) = 8 + 24 - 8 - 24 = 0 (by factor theorem)

OpenStudy (precal):